OpenStudy (kc_kennylau):

Identify the mistake in this "proof": (1 point) \[\Large\begin{array}{rcl&r} 0&=&0^1&\mbox{step 1}\\ &=&0^{2-1}&\mbox{step 2}\\ &=&\frac{0^2}{0^1}&\mbox{step 3}\\ &=&\frac00&\mbox{step 4}\\ &=&\text{undefined}&\mbox{step 5} \end{array}\\ \therefore\mbox{0 is undefined.}\]

OpenStudy (anonymous):

hmm... i asked a question simiar to this I've asked before I think I can help you

OpenStudy (kc_kennylau):

@Rancd @HELP-ME-PLZ

OpenStudy (kc_kennylau):

lolz using tags to write a msg

OpenStudy (anonymous):

The 2 part is wrong though but I need to prove why

OpenStudy (anonymous):

I know why

OpenStudy (anonymous):

Thats not how you prove you need to multiply not divde For let's say x is a number any numbe ror value x * 0 = 0/0 * 0

OpenStudy (anonymous):

Hence x can be any number that can go into 0 but how do we prove if its undefined

OpenStudy (anonymous):

step 1 and 2 are true and have real solutions, step three has no solution and is extraneous

OpenStudy (kc_kennylau):

but what is wrong with the transition from step 2 to step 3?

OpenStudy (anonymous):

because the two are not the same, you can't divide by 0

OpenStudy (kc_kennylau):

I'm asking why are those 2 not the same

OpenStudy (kc_kennylau):

Which rule has it violated to change from step 2 to step 3

OpenStudy (anonymous):

1 =2

OpenStudy (anonymous):

because in the case that you are using it, it is undefined. 0^2 is the same as 0 times 0 and it = 0. it is like you are trying to divide 0^2 by 0 to prove it, and you can't do that.

OpenStudy (kc_kennylau):

This is the answer: Step 3 is invalid, as the identity \(\Large a^{m-n}=\frac{a^m}{a^n}\) is only valid IF a is not 0, and m>n>0. *sigh* you guys...

OpenStudy (anonymous):

what do you do this to us ?

OpenStudy (anonymous):

Give us another question ?

OpenStudy (kc_kennylau):

I created this question xpp

OpenStudy (anonymous):

oh, you wanted the technical answer, well sheesh

OpenStudy (anonymous):

I DEMAND another one por favoe !!

OpenStudy (kc_kennylau):

uh, i need to think...

OpenStudy (kc_kennylau):

so hard to think

OpenStudy (anonymous):

Never I'm getting revenge instead

OpenStudy (anonymous):

https://brilliant.org/mathematics-problem/looking-for-integers/?group=EcAAcZbU1fIj Solve this you've got 5 minutes

OpenStudy (anonymous):

Bye Ciao

OpenStudy (kc_kennylau):

wow so easy, took me 2 trials

OpenStudy (anonymous):

You saw a shirt for $97 You didn't have the money so you borrowed $50 from your mom and $50 from your dad. $50 + $50 = $100 You bought the shirt and had $3 change. You gave yor dad $1 and your mom $1 and kept the other $1 to yourself. Now you owe your mom $49 and your dad $49. $49 + $49 = $98 + your $1 = $99 Where is the missing $1 ?

OpenStudy (kc_kennylau):

It makes no sense adding the money you owe your mom and your dad and the money you currently have together

OpenStudy (kc_kennylau):

I have to go now, see you guys, have a happy new year

OpenStudy (anonymous):

Happy new year In 5 days I will create a question even you can't solve